write the equation of the line with a slope of 3/7 passing through the points (9, -2)
Answer by tutorcecilia(2152)

First Method:
(y-y2)=m(x-x2) [Use the point-slope format]
y-(-2)=(3/7)(x-9) [Plug-in the given values: m=(3/7) and points-(9, -2)
y+2=3/7x-(27/7) [Solve for y]
y=(3/7)x-(27/7)-2
y=(3/7)x-(41/7) [Final answer]
.
.
Second Method:
y=mx+b [Use the slope-intercept format]
(-2)=(3/7)(9)+b [Plug-in the given values: m=(3/7) and points-(9, -2)
-2=(27/7) + b [Solve for "b"]
(-14-27)/7 = b [Simplify]
-41/7 = b
y=(3/7)x-(41/7) [Plug-in the values for "m" and "b"][final answer]